OpenStudy (anonymous):

Laser light with an intensity of 38300 W/m2 perpendicularly hits a black perfect absorber rectangular area of width W = 10 mm and height H = 10 mm. What is the force (in N) this laser exerts on the area?

OpenStudy (anonymous):

Any help would be greatly appreciated!!

OpenStudy (anonymous):

light intensity equation is \[I = \frac{ Power }{ Area } = \frac{ F.V }{ A } .\] Thus\[F = \frac{ I A }{ V }.\] and the velocity of light is constant.

OpenStudy (anonymous):

@Catch.me I have tried doing that and it tells me I am wrong 38300*(.01*.01)/3E8=1.28E8

OpenStudy (anonymous):

Just to be clear, do you mean 1.28E-8 ?

OpenStudy (anonymous):

@witthan1 it's power -8 as @Pompeii00 said.